OpenStudy (anonymous):

How many atoms are in 1.14 mol so3?

OpenStudy (anonymous):

1 mole of SO3 will contain 6.02 x 10^23 mol^-1 of SO3 molecules. thus, 1.14moles will contain; = 1.14mol x [3mol O/1mol SO3] x [6.02 x 10^23 atoms O/1mol O] = 2.05884 x 10^24 oxygen atoms = 1.14mol x [1mol S/1mol SO3] x [6.02 x 10^23 atoms O/1mol O] = 6.8628 x 10^23 sulfur atoms

OpenStudy (anonymous):

lol yahoo answers!

OpenStudy (anonymous):

living the dream

OpenStudy (anonymous):

looooool!!!